A tank holds 80 liters of chemical solution. Currently, the solution has a strength of 30% How much of this solution must be drained and replace with a 70% solution to get a strength of 40%
Answer by Boreal(15235)

amount drained and replaced=x
(80-x)0.30x+.70x=(80)*0.4
24-.3x+.70x=32
.4x+24=32
.4x=8
x=20 liters
Check
Start with 80*.30=24 liters of pure.
Want 80 liters of .40=32 liters of pure
For every 10 liters of 30% removed and replaced with 70%, 4 extra liters of pure are added.
Need 20 liters of 30% removed and replaced with 70%.
That will be 60 liters*.30=18 liters and 20 liters of 70% = 14 liters. Total 32 liters.